# Overcoming False Illusions in Real-World Face Restoration with Multi-Modal Guided Diffusion Model

### v1

@@ Abstract
We introduce a novel Multi-modal Guided Real-World Face Restoration (MGFR) technique designed to improve the quality of facial image restoration from low-quality inputs. Leveraging a blend of attribute text prompts, high-quality reference images, and identity information, MGFR can mitigate the generation of false facial attributes and identities often associated with generative face restoration methods. By incorporating a dual-control adapter and a two-stage training strategy, our method effectively utilizes multi-modal prior information for targeted restoration tasks. We also present the Reface-HQ dataset, comprising over 23,000 high-resolution facial images across 5,000 identities, to address the need for reference face training images. Our approach achieves superior visual quality in restoring facial details under severe degradation and allows for controlled restoration processes, enhancing the accuracy of identity preservation and attribute correction. Including negative quality samples and attribute prompts in the training further refines the model’s ability to generate detailed and perceptually accurate images.
@@ 1 Introduction
Real-World Face Restoration (FR) aims to reconstruct high-resolution, high-quality (HQ) facial images from their degraded, low-resolution observations. Recent works, leveraging powerful generative priors and diffusion models, have achieved significant progress Menon et al. (2020); Yang et al. (2021b); Wang et al. (2021b); Lin et al. (2023); Wang et al. (2023b), particularly in addressing severely degraded facial images. However, the information contained in the low-quality (LQ) inputs is limited. FR inevitably introduces the illusion of generation, producing results with different facial attributes or even different identities from the target image. For example, in Figure 1 (a) and Figure 2, we cannot effectively predict the eye colour and skin characteristics of the person in the LQ input, resulting in the output results – even the quality can be improved – having an apparent perceptual distance from the target image. Many applications find this unacceptable, as humans can readily identify these flaws. Achieving optimal facial image recovery requires effectively tackling false hallucinations.
@@
This work proposes a method called Multi-modal Guided Real-World Face Restoration (MGFR). We aim to use multiple control methods to consider diverse multi-modal prior information in FR to restore face images in a targeted manner. Specifically, MGFR uses attribute text prompts, HQ reference images, and identity information as priors for collaborative guidance during restoration. We design a dual-control adapter with a two-stage training strategy to balance the complex multi-modal and multi-source prior information. This dual controller is compatible with pre-trained generative diffusion models Rombach et al. (2022) and prioritises restoration tasks while incorporating additional multi-modal guidance. In addition, we collect the Reface-HQ dataset to address the scarcity of reference image samples containing over 5000 identities and 23000 high-resolution facial images. Based on the FFHQ Karras et al. (2019a) and the proposed Reface-HQ datasets, we develop a high-quality synthetic dataset for model training enriched with attribute text prompts. Furthermore, we adopt a counter-intuitive strategy to integrate negative-quality samples with negative-quality prompts and negative-attribute prompts into training to enhance perceptual quality and detail generation.
@@
Figure 2: Motivation. In conditions of severe degradation, the loss of facial identity information becomes profoundly pronounced without reference image. During the face restoration process, distortions of facial attributes, including gender and age, are commonly encountered. Appropriate attribute prompts can offer additional reference points and exert control in the recovery process.
@@ 2 Related Works
Real-World face restoration (FR) concentrates on the challenging task of reconstructing HQ face images from LQ inputs. These LQ inputs are often blemished by various forms of quality degradation such as low-resolution Chen et al. (2018); Dong et al. (2014); Lim et al. (2017), blur Kupyn et al. (2018); Shen et al. (2018), noise Zhang et al. (2017), and JPEG compression artifacts Dong et al. (2015), etc. FR heavily relies on facial priors, such as facial landmarks Chen et al. (2018), parsing maps Chen et al. (2018; 2021), and facial component heatmaps Yu et al. (2018). Generative priors Karras et al. (2020); Rombach et al. (2022); Gu et al. (2020); Shen et al. (2020) have also emerged as fundamental elements in providing vibrant textures and details in FR Menon et al. (2020); Hu et al. (2023); Zhu et al. (2022). Advanced techniques like GPEN Yang et al. (2021a), GFP-GAN Wang et al. (2021a), and GLEAN Chan et al. (2021) are recognized for more effectively incorporating these priors within encoder-decoder structures. There are also works that considerably reduce the uncertainty commonly associated with generative priors Gu et al. (2022); Zhou et al. (2022); Wang et al. (2022), which are trained on discrete feature codebooks for high-quality facial images. Recently, diffusion models like DiffBIR Lin et al. (2023) have revitalized interest in this area, leveraging the generative power of pre-trained LDM as a prior. DR2 Wang et al. (2023b) also makes contributions by transforming input images into noisy states and then iteratively denoising them to capture the essential semantic information.
@@
Reference-based face restoration utilizes HQ images of the same identity as references. This concept was first introduced in Li et al. (2018a). To address discrepancies in poses and expressions, GWAInet Dogan et al. (2019) and the later work of Li et al. Li et al. (2020; 2018b) focused on more effectively directing deformations or choosing the optimal reference image for reconstruction. MyStyle Nitzan et al. (2022) adopts a unique approach by refining StyleGAN Karras et al. (2019a) with numerous reference images based on personal appearance. DMDNet Li et al. (2023) employs a dictionary constructed from diverse, high-quality facial images to rehabilitate degraded images using its high-quality components. In the MGFR framework, incorporating a single reference image is vital for tailoring the restoration to individual faces. Unlike conventional methods, MGFR does not require strict alignment constraints on expressions or postures.
@@
Diffusion models have shown significant effectiveness in a broad range of image processing tasks. Current methods Chen et al. (2023b); Zhang et al. ; Yu et al. (2024); Chen et al. (2023a) employ pre-trained text-to-image diffusion models Rombach et al. (2022) for image processing, demonstrating the potential of language as a comprehensive input for image reconstruction tasks. Concurrently, approaches like ControlNet Zhang et al. , T2I-adapter Mou et al. (2023), and ControlNet-XS Zavadski et al. (2023) have further developed the integration of more intricate condition controls within the text-to-image framework, facilitating more precise and tailored image generation. Nevertheless, the field of FR, particularly in the utilization of natural language prompts, continues to be an area of untapped potential.
@@ 3.1 Stage One: Text-guided Face Restoration
This stage trains the face image restoration model that accepts text prompt as input, as shown in Figure 5. We use the pre-trained Stable Diffusion (SD) Rombach et al. (2022) model as our generative prior and train an additional adapter to extend it to the face restoration applications. The pre-trained SD generative prior has the ability to understand face image attributes text and generate high-quality face images. In this stage, our model restores the image $x$ according to the condition $\{y,c_{a}\}$, where $y$ represents the degraded LQ image, and $c_{a}$ constitutes the facial attribute prompts describing the face attributes. We first use the CLIP text encoder Radford et al. (2021) to calculate the text embedding $e_{r}=\operatorname{CLIP}(c_{a})$. The LQ input $y$ is also mapped to a latent representation $z_{lq}$ using the VAE encoder in SD Rombach et al. (2022). We then perform diffusion generation on this latent representation. In the framework of SD, the model uses UNet Ronneberger et al. (2015) denoising model ${\mathcal{E}}_{\theta}\left({z}_{t},t,e_{r}\right)$ to perform the diffusion generation process, where $t$ is the time stamp in diffusion model and $z_{t}$ is the intermediate results at time $t$. Based on the ControlNet Zhang et al. framework, we use an external adapter that takes the LQ input $y$ and text prompts embedding $e_{r}$ as input to provide guidance for the fixed UNet $\mathcal{E}_{\theta}$. We call this adapter the LQ Control Adapter (LCA). Specifically, the UNet model contains the encoder, intermediate blocks, and the decoder. The decoder receives features from the encoder and fuses them at each corresponding scale. The LCA contains the same encoder and intermediate blocks as in the UNet model. The feature output of each scale in LCA is integrated with the corresponding scale of the UNet decoder to achieve the effect of output control. However, we found that simply using the above ControlNet framework has a key limitation – the lack of information exchange from the UNet encoder to the LCA. This gap means that the LCA is unaware of the processes that are performed in the UNet encoder, thereby limiting its ability to generate effective control features. In order to solve this problem, we add the feature output of each scale in the UNet encoder to the corresponding scale in the LCA. The LQ controller part of Figure 5 illustrates this operation. In this way, the capability of the LCA is greatly enhanced, so better visual effects and control results can be achieved.
@@ 3.2 Stage Two: Multi-modal-guided Face Restoration
After the first stage of training, the model can already reconstruct high-quality images from the LQ inputs, guided by text prompts. Next, we further enrich the guidance and introduce high-quality reference images and face identity information as additional control means based on the first-stage model. We design a new Dual-Control Adapter (DCA), as shown in Figure 3. In DCA, we introduce a Reference Control Adapter (RCA) specifically for reference image processing. RCA has the same architecture as LCA, and its role is to extract related and useful information from reference images and identity information and provide additional details to LCA. The input of RCA includes an HQ reference face image $r$ containing the same identity as the LQ input and its identity information embedding $e_{f}$. For the reference image $r$, we first use the VAE encoder consistent with the SD model for feature extraction to obtain $z_{ref}$. Next, we fuse the LQ latent representation $z_{lq}$ with $z_{ref}$ using a reference and LQ feature fusion module (RLF). This module allows RCA to identify the high-frequency details missing in the LQ input and perform targeted information extraction for restoration guidance. For identity information embedding, we calculate $e_{f}=\operatorname{Proj}(\operatorname{Arcface}(r))$, where $\operatorname{Arcface}(r)$ is the face recognition model Arcface Deng et al. (2019) to extract the identity feature from the reference image $r$. We align it to the space that RCA can handle through a trainable linear projection layer $\operatorname{Proj}(\cdot)$. Due to the function of the RCA extracting information from the reference image according to the LQ input, the RCA requires the information of the LCA branch as input. At the same time, RCA needs to provide the extracted information back to LCA in reverse. Therefore, we designed a dual-way interaction mechanism for RCA and LCA, as shown in Figure 5. In this design, RCA runs in parallel with LCA. At each scale, the LQ block in LCA first processes the fused information of both two branches and then hands the intermediate features to RCA. RCA performs feature extraction and processing based on these intermediate results and reference conditions and finally uses the same operation to apply the processing results to the next layer of LCA processing. Finally, the output of each scale of LCA is applied to the corresponding position of the UNet decoder. RCA directly affects LCA and, therefore, also affects the calculation of UNet. Since then, we have had a dual-control adapter that can accept multiple control inputs.
@@ 3.3 Negative Samples and Prompt
Classifier-Free Guidance (CFG) Ho & Salimans (2022) introduces a novel control mechanism utilizing negative prompts to delineate unwanted content for the model. This feature can be leveraged to inhibit the generation of low-quality images by the model and to enhance the precision of facial detail reconstruction. Throughout the inference phase, at each step of diffusion, three distinct predictions are generated: one employing the positive prompt $pos$, another using the negative quality prompt $nq$, and a third via the negative attribute prompt $na$ (the negation sentence described by $pos$). We combine the results generated from these different prompts to form the final output:
@@ 4.1 Experimental Setting
Datasets. Our two-stage training method requires different datasets for training. For the first stage, we mainly train the model’s ability to restore HQ images and process text prompts. Therefore, we need HQ images with text annotations for training. We synthesize training image pairs using the FFHQ dataset Karras et al. (2019b). FFHQ contains 70,000 high-resolution face images, and we resize these images to $512\times 512$ for training. In the second stage, in addition to requiring HQ face images to create training image pairs, we also need to assign HQ reference images with consistent identities but different details to each image. Although there are some datasets proposed for reference face restoration Liu et al. (2015); Yi et al. (2014), the resolution and quality of these datasets cannot meet the current requirements. In this work, we collect a new dataset for referenced face restoration called Reface-HQ. Reface-HQ contains 23,500 high-quality and diverse images of over 5,250 identities. Additional details of Reface-HQ can be found in the Appendix A. To synthesize LQ images, we follow the degradation model and setting used in Wang et al. (2023b). Our test data also involves multiple sources, including CelebA-Test Liu et al. (2015), Reface-Test, and real-world LQ images collected from the Internet. Specifically, CelebA-Test contains 3,000 testing images from the CelebA-HQ dataset. Reface-Test contains 1,800 images of 380 identities split from the proposed Reface-HQ dataset. The LQ images for testing are synthesized within the same degradation range as the training setting.
@@
Figure 9: Training Data Composition. Initially, negative quality samples are incorporated into the training to enhance the clarity and quality of the restored image. Furthermore, large language models coupled with a facial attribute classifier are employed to extract attribute texts for integration into the training.
@@
Attribute Prompt. Text prompts are important for us to control face attributes and improve quality. In our method, a total of three types of prompts are introduced. Two attribute prompts describe the face attributes, and the last one describes the negative quality of the image. For attribute prompts, $pos$ contains positive descriptions of face attributes, while $na$ describes attributes that do not exist in this image to provide negative prompts of attributes. To obtain these descriptions, we first use a pre-trained face attribute detector He et al. (2017) to extract the presence of each attribute in the face. We considered 28 different attributes in this work. For attributes that have high confidence to exist, we add them to the $pos$ positive attributes. For the remaining attributes with low confidence, we classify them as $na$ negative attributes. At this time, these attributes are still separate words. We use a large language model to organize the separated words into natural language to facilitate the understanding of the CLIP text encoder. Thus, each face image is associated with two attribute prompts detailing existing and non-existing attributes. For the negative quality prompt, $nq$ involves “low quality, low resolution, over-smoothed and distorted images”, as shown in Figure 9 (a). See Appendix B for comprehensive details on the training and inference procedures involving attribute prompts.
@@
Implementation. The training involved fine-tuning based on Stable Diffusion v2.1 Rombach et al. (2022), with the control adapter structure adhering to Zhang et al. . The Adam optimizer Kingma & Ba (2014) was employed, featuring a learning rate of $e^{-5}$. The initial training stage spanned 15 days, while the subsequent stage lasted 5 days, utilizing 4 Nvidia A100 GPUs with a batch size of 4. For testing purposes, the hyperparameters were set as $T=500$, $\lambda_{na}$ = 0.5 and $\lambda_{nq}$ = 0.5.
@@
Metrics. For quantitative comparison, followed by many previous works Lin et al. (2023); Yu et al. (2024), the selected metrics include full-reference metrics PSNR, SSIM, and LPIPS Zhang et al. (2018), as well as non-reference metrics ManIQA Yang et al. (2022), ClipIQA Wang et al. (2023a), and MUSIQ Ke et al. (2021). Furthermore, the Arcface identity distance Deng et al. (2019) (ID) is utilized to assess the similarity of identity information.

### v2

@@ Abstract
We introduce a novel Multi-modal Guided Real-World Face Restoration (MGFR) technique designed to improve the quality of facial image restoration from low-quality inputs. Leveraging a blend of attribute text prompts, high-quality reference images, and identity information, MGFR can mitigate the generation of false facial attributes and identities often associated with generative face restoration methods. By incorporating a dual-control adapter and a two-stage training strategy, our method effectively utilizes multi-modal prior information for targeted restoration tasks. We also present the Reface-HQ dataset, comprising over 21,000 high-resolution facial images across 4800 identities, to address the need for reference face training images. Our approach achieves superior visual quality in restoring facial details under severe degradation and allows for controlled restoration processes, enhancing the accuracy of identity preservation and attribute correction. Including negative quality samples and attribute prompts in the training further refines the model’s ability to generate detailed and perceptually accurate images.
@@ 1 Introduction
Real-World Face Restoration (FR) aims to reconstruct high-resolution, high-quality (HQ) facial images from their degraded, low-resolution observations. Recent works, leveraging powerful generative priors and diffusion models, have achieved significant progress (Menon et al., 2020; Yang et al., 2021b; Wang et al., 2021b; Lin et al., 2023; Wang et al., 2023b), particularly in addressing severely degraded facial images. However, the information contained in the low-quality (LQ) inputs is limited. FR inevitably introduces the illusion of generation, producing results with different facial attributes or even different identities from the target image. For example, in Figure 1 (a) and Figure 2, we cannot effectively predict the eye colour and skin characteristics of the person in the LQ input, resulting in the output results – even the quality can be improved – having an apparent perceptual distance from the target image. Many applications find this unacceptable, as humans can readily identify these flaws. Achieving optimal facial image recovery requires effectively tackling false hallucinations.
@@
This work proposes a method called Multi-modal Guided Real-World Face Restoration (MGFR). We aim to use multiple control methods to consider diverse multi-modal prior information in FR to restore face images in a targeted manner. Specifically, MGFR uses attribute text prompts, HQ reference images, and identity information as priors for collaborative guidance during restoration. We designed a dual-control adapter with a two-stage training strategy to balance the complex multi-modal and multi-source prior information. This dual controller is compatible with pre-trained generative diffusion models (Rombach et al., 2022) and prioritizes restoration tasks while incorporating additional multi-modal guidance. In addition, we collect the Reface-HQ dataset to address the scarcity of reference image samples containing over 4800 identities and 21000 high-resolution facial images. Based on the FFHQ (Karras et al., 2019a) and the proposed Reface-HQ datasets, we develop a high-quality synthetic dataset for model training enriched with attribute text prompts. Furthermore, we adopt a counterintuitive strategy to integrate negative-quality samples with negative-quality prompts and negative-attribute prompts into training to enhance perceptual quality and detail generation.
@@
Figure 2: Motivation. In conditions of severe degradation, the loss of facial identity information becomes profoundly pronounced without a reference image. During the face restoration process, distortions of facial attributes, including gender and age, are commonly encountered. Appropriate attribute prompts can offer additional reference points and exert control in the recovery process.
@@ 2 Related Works
Real-world face restoration (FR) concentrates on the challenging task of reconstructing HQ face images from LQ inputs. These LQ inputs are often blemished by various forms of quality degradation, such as low-resolution (Chen et al., 2018; Dong et al., 2014; Lim et al., 2017), blur (Kupyn et al., 2018; Shen et al., 2018), noise (Zhang et al., 2017), and JPEG compression artifacts (Dong et al., 2015), etc. FR heavily relies on facial priors, such as facial landmarks (Chen et al., 2018), parsing maps (Chen et al., 2018; 2021), and facial component heatmaps (Yu et al., 2018). Generative priors (Karras et al., 2020; Rombach et al., 2022; Gu et al., 2020; Shen et al., 2020) have also emerged as fundamental elements in providing vibrant textures and details in FR (Menon et al., 2020; Hu et al., 2023; Zhu et al., 2022). Advanced techniques like GPEN (Yang et al., 2021a), GFP-GAN (Wang et al., 2021a), and GLEAN (Chan et al., 2021) are recognized for more effectively incorporating these priors within encoder-decoder structures. There are also works that considerably reduce the uncertainty commonly associated with generative priors (Gu et al., 2022; Zhou et al., 2022; Wang et al., 2022), which are trained on discrete feature codebooks for high-quality facial images. Recently, diffusion models like DiffBIR (Lin et al., 2023) have revitalized interest in this area, leveraging the generative power of pre-trained LDM as a prior. DR2 (Wang et al., 2023b) also contributes by transforming input images into noisy states and then denoising them to capture the essential semantic information.
@@
Reference-based face restoration utilizes HQ images of the same identity as references. This concept was first introduced in (Li et al., 2018a). To address discrepancies in poses and expressions, GWAInet (Dogan et al., 2019) and the later work of Li et al. (Li et al., 2020b; 2018b) focused on more effectively directing deformations or choosing the optimal reference image for reconstruction. MyStyle (Nitzan et al., 2022) adopts a unique approach by refining StyleGAN (Karras et al., 2019a) with numerous reference images based on personal appearance. DMDNet (Li et al., 2023) employs a dictionary constructed from diverse, high-quality facial images to rehabilitate degraded images using its high-quality components. In the MGFR framework, incorporating a single reference image is vital for tailoring the restoration to individual faces. Unlike conventional methods, MGFR does not require strict alignment constraints on expressions or postures.
@@
Diffusion models have shown significant effectiveness in a broad range of image processing tasks. Current methods (Chen et al., 2023b; Zhang et al., 2023; Yu et al., 2024; Chen et al., 2023a) employ pre-trained text-to-image diffusion models (Rombach et al., 2022) for image processing, demonstrating the potential of language as a comprehensive input for image reconstruction tasks. Concurrently, approaches like ControlNet (Zhang et al., 2023), T2I-adapter (Mou et al., 2023), and ControlNet-XS (Zavadski et al., 2023) have further developed the integration of more intricate condition controls within the text-to-image framework, facilitating more precise and tailored image generation. Nevertheless, the field of FR, particularly in the utilization of natural language prompts, continues to be an area of untapped potential.
@@ 3.1 Stage One: Text-guided Face Restoration
This stage trains the face image restoration model that accepts text prompt as input, as shown in Figure 5. We use the pre-trained Stable Diffusion (SD) (Rombach et al., 2022) model as our generative prior and train an additional adapter to extend it to the face restoration applications. The pre-trained SD generative prior has the ability to understand face image attributes and text and generate high-quality face images. In this stage, our model restores the image $x$ according to the condition $\{y,c_{a}\}$, where $y$ represents the degraded LQ image, and $c_{a}$ constitutes the facial attribute prompts describing the face attributes. We first use the CLIP text encoder (Radford et al., 2021) to calculate the text embedding $e_{r}=\operatorname{CLIP}(c_{a})$. The LQ input $y$ is also mapped to a latent representation $z_{lq}$ using the VAE encoder in SD (Rombach et al., 2022). We then perform diffusion generation on this latent representation. In the framework of SD, the model uses UNet (Ronneberger et al., 2015) denoising model ${\mathcal{E}}_{\theta}\left({z}_{t},t,e_{r}\right)$ to perform the diffusion generation process, where $t$ is the time stamp in diffusion model and $z_{t}$ is the intermediate results at time $t$. Based on the ControlNet (Zhang et al., 2023) framework, we use an external adapter that takes the LQ input $y$ and text prompts embedding $e_{r}$ as input to provide guidance for the fixed UNet $\mathcal{E}_{\theta}$. We call this adapter the LQ Control Adapter (LCA). Specifically, the UNet model contains the encoder, intermediate blocks, and the decoder. The decoder receives features from the encoder and fuses them at each corresponding scale. The LCA contains the same encoder and intermediate blocks as the UNet model. The feature output of each scale in LCA is integrated with the corresponding scale of the UNet decoder to achieve the effect of output control. However, we found that simply using the above ControlNet framework has a key limitation – the lack of information exchange from the UNet encoder to the LCA. This gap means that the LCA is unaware of the processes that are performed in the UNet encoder, thereby limiting its ability to generate effective control features. In order to solve this problem, we add the feature output of each scale in the UNet encoder to the corresponding scale in the LCA. The LQ controller part of Figure 5 illustrates this operation. In this way, the capability of the LCA is greatly enhanced, so better visual effects and control results can be achieved.
@@ 3.2 Stage Two: Multi-modal-guided Face Restoration
After the first stage of training, the model can already reconstruct high-quality images from the LQ inputs guided by text prompts. Next, we further enrich the guidance and introduce high-quality reference images and face identity information as additional control means based on the first-stage model. We design a new Dual-Control Adapter (DCA), as shown in Figure 3. In DCA, we introduce a Reference Control Adapter (RCA) specifically for reference image processing. RCA has the same architecture as LCA, and its role is to extract related and useful information from reference images and identity information and provide additional details to LCA. The input of RCA includes an HQ reference face image $r$ containing the same identity as the LQ input and its identity information embedding $e_{f}$. For the reference image $r$, we first use the VAE encoder consistent with the SD model for feature extraction to obtain $z_{ref}$. Next, we fuse the LQ latent representation $z_{lq}$ with $z_{ref}$ using a reference and LQ feature fusion module (RLF). This module allows RCA to identify the high-frequency details missing in the LQ input and perform targeted information extraction for restoration guidance. For identity information embedding, we calculate $e_{f}=\operatorname{Proj}(\operatorname{Arcface}(r))$, where $\operatorname{Arcface}(r)$ is the face recognition model Arcface (Deng et al., 2019) to extract the identity feature from the reference image $r$. We align it to the space that RCA can handle through a trainable linear projection layer $\operatorname{Proj}(\cdot)$. Due to the function of the RCA extracting information from the reference image according to the LQ input, the RCA requires the information of the LCA branch as input. At the same time, RCA needs to provide the extracted information back to LCA in reverse. Therefore, we designed a dual-way interaction mechanism for RCA and LCA, as shown in Figure 5. In this design, RCA runs in parallel with LCA. At each scale, the LQ block in LCA first processes the fused information of both two branches and then hands the intermediate features to RCA. RCA performs feature extraction and processing based on these intermediate results and reference conditions and finally uses the same operation to apply the processing results to the next layer of LCA processing. Finally, the output of each scale of LCA is applied to the corresponding position of the UNet decoder. RCA directly affects LCA and, therefore, also affects the calculation of UNet. Since then, we have had a dual-control adapter that can accept multiple control inputs.
@@ 3.3 Negative Samples and Prompt
Classifier-Free Guidance (CFG) (Ho & Salimans, 2022) introduces a novel control mechanism utilizing negative prompts to delineate unwanted content for the model. This feature can be leveraged to inhibit the generation of low-quality images by the model and to enhance the precision of facial detail reconstruction. Throughout the inference phase, at each step of diffusion, three distinct predictions are generated: one employing the positive prompt $pos$, another using the negative quality prompt $nq$, and a third via the negative attribute prompt $na$ (the negation sentence described by $pos$). We combine the results generated from these different prompts to form the final output:
@@ 4.1 Experimental Setting
Datasets. Our two-stage training method requires different datasets for training. For the first stage, we mainly train the model’s ability to restore HQ images and process text prompts. Therefore, we need HQ images with text annotations for training. We synthesize training image pairs using the FFHQ dataset (Karras et al., 2019b). FFHQ contains 70,000 high-resolution face images, and we resize these images to $512\times 512$ for training. In the second stage, in addition to requiring HQ face images to create training image pairs, we also need to assign HQ reference images with consistent identities but different details to each image. Although there are some datasets proposed for reference face restoration (Liu et al., 2015; Yi et al., 2014), the resolution and quality of these datasets cannot meet the current requirements. In this work, we collect a new dataset for referenced face restoration called Reface-HQ. Reface-HQ contains 21,500 high-quality and diverse images of over 4800 identities. Additional details of Reface-HQ can be found in the Appendix A. To synthesize LQ images, we follow the degradation model and setting used in (Wang et al., 2023b). Our test data also involves multiple sources, including CelebA-Test (Liu et al., 2015), Reface-Test, and real-world LQ images collected from the Internet. Specifically, CelebA-Test contains 3,000 testing images from the CelebA-HQ dataset. Reface-Test contains 1,300 images of 280 identities split from the proposed Reface-HQ dataset. The LQ images for testing are synthesized within the same degradation range as the training setting.
@@
Figure 9: Training Data Composition. Initially, negative quality samples are incorporated into the training to enhance the clarity and quality of the restored image. Furthermore, large language models, coupled with a facial attribute classifier, are employed to extract attribute texts for integration into the training.
@@
Attribute Prompt. Text prompts are important for us to control face attributes and improve quality. In our method, a total of three types of prompts are introduced. Two attribute prompts describe the face attributes, and the last one describes the negative quality of the image. For attribute prompts, $pos$ contains positive descriptions of face attributes, while $na$ describes attributes that do not exist in this image to provide negative prompts of attributes. To obtain these descriptions, we first use a pre-trained face attribute detector (He et al., 2017) to extract the presence of each attribute in the face. We considered 28 different attributes in this work. For attributes that have high confidence to exist, we add them to the $pos$ positive attributes. For the remaining attributes with low confidence, we classify them as $na$ negative attributes. At this time, these attributes are still separate words. We use a large language model to organize the separated words into natural language to facilitate the understanding of the CLIP text encoder. Thus, each face image is associated with two attribute prompts detailing existing and non-existing attributes. For the negative quality prompt, $nq$ involves “low quality, low resolution, over-smoothed and distorted images”, as shown in Figure 9 (a). See Appendix B for comprehensive details on the training and inference procedures involving attribute prompts.
@@
Implementation. The training involved fine-tuning based on Stable Diffusion v2.1 (Rombach et al., 2022), with the control adapter structure adhering to (Zhang et al., 2023). The Adam optimizer (Kingma & Ba, 2014) was employed, featuring a learning rate of $e^{-5}$. The initial training stage spanned 15 days, while the subsequent stage lasted 5 days, utilizing 4 Nvidia A100 GPUs with a batch size of 4. For testing purposes, the hyperparameters were set as $T=500$, $\lambda_{na}$ = 0.5 and $\lambda_{nq}$ = 0.5.
@@
Metrics. For quantitative comparison, followed by many previous works (Lin et al., 2023; Yu et al., 2024), the selected metrics include full-reference metrics PSNR, SSIM, and LPIPS (Zhang et al., 2018), as well as non-reference metrics ManIQA (Yang et al., 2022), ClipIQA (Wang et al., 2023a), and MUSIQ (Ke et al., 2021). Furthermore, the Arcface identity distance (Deng et al., 2019) (ID) is utilized to assess the similarity of identity information.